OpenStudy (anonymous):

Jim is 2 years older than 3 times his little brother,tommy.together their ages add up to 18.how old is each?

OpenStudy (anonymous):

Steps: 1) 2 + 3x = 18

OpenStudy (anonymous):

2) Jim 2+3x =18 3x=18-2 3x=16 x=5.3333

OpenStudy (anonymous):

3) 2+3(5.3) = 18

OpenStudy (anonymous):

Kid style

OpenStudy (anonymous):

????

OpenStudy (anonymous):

i dont understand you

OpenStudy (anonymous):

Easy style I'm in 6 grade

OpenStudy (anonymous):

ok where did i loose you

OpenStudy (anonymous):

5.3

OpenStudy (anonymous):

when you divide 16 by 3 you get 5.3333 which is the same as 5.3. You then insert that into the equation.

OpenStudy (anonymous):

3x5.3=? The answer you get you subtract from 2 in order to get little brother. THen to find Jim add 2 years to tommy's age

OpenStudy (anonymous):

is this any simpler?